A light aircraft of mass 1200 kg starts from rest and accelerates along a straight horizontal runway. Their aircraft engine produces a constant thrust of 3400 Newtons.A constant friction force of 400 Newton acts on the aircraft.the aircraft takes off when it reaches a speed of 35 metres per second.calculate the distance travelled by an aircraft from a standing point until it takes off.
1
Expert's answer
2020-12-17T09:03:29-0500

Calculate the acceleration by Newton's second law:


"a=\frac{F_\text{net}}{m}=\frac{F-f}{m}."


Taking into account the plane started from rest, the distance it traveled is


"d=\frac{v^2}{2a}=\frac{mv^2}{2(F-f)},\\\space\\
d=\frac{1200\cdot35^2}{2(3400-400)}=245\text{ m}."